OpenStudy (anonymous):

There are nickels and dimes in a candy jar. If there are 22 more nickels than dimes write an expression that represents of the total number of coins if the number of dimes is d.

OpenStudy (anonymous):

n=d+22 d+n=d+(d+22) so the equation should be something like: (total amount of coins= d+(d+22) )

OpenStudy (anonymous):

d+(d+22) ? are you sure..

OpenStudy (anonymous):

@ineedhelpnow08

OpenStudy (anonymous):

He's definitely right.

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

now people need to help me